OpenStudy (anonymous):

The sum of two real numbers is 14 and the product is 38. What are the two numbers?

OpenStudy (mathmate):

Let one of the two numbers be x, then the other number is 14-x. If the product is 38, it means that when multiplied together, the result is 38, or \[x (14-x)=38\] Expand and solve by the quadratic formula to find (two) values of x. \[14x-x^2-38=0\]

OpenStudy (anonymous):

7 plus or minus sqaure root of 11@mathmate

OpenStudy (anonymous):

The diagonal of a square measures 6 inches. What is the length of the sides?

OpenStudy (mathmate):

Call the side s. Then the \[diagonal =\sqrt{s^2+s^2}=\sqrt{2}\ s \] Equate to 6 inches and solve for s.

OpenStudy (anonymous):

1 @mathmate

OpenStudy (mathmate):

By the way, your answer for the first problem (7 +/- sqrt(11) ) is correct. Well done!

OpenStudy (anonymous):

6 square of 2 @mathmate

OpenStudy (anonymous):

A new computer can complete a complex biochemistry calculation in 11 hours less than an older-model computer. Working together, the new computer and the old computer can complete the calculation in 19 hours. How long would it take the new computer to complete the calculation by itself? Round to the nearest tenth of an hour. Select one: a. 33.3 hours b. 19.0 hours c. 24.3 hours d. 28.8 hours

OpenStudy (mathmate):

\[\sqrt{2}*s = 6\] so \[s=6\div\sqrt{2} =3\sqrt{2}\]

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

The owner of a hobby store bought a case of 9-volt batteries for $47.80. He marked the price up such that his profit was $3.55 per battery. Once he had sold 12 batteries, the store owner had recovered his cost. how many 9-volt batteries are in a case? Select one: a. 120 batteries b. 105 batteries c. 110 batteries d. 115 batteries @mathmate

OpenStudy (mathmate):

Do you still want the computer problem done?

OpenStudy (anonymous):

yes please . @mathmate

OpenStudy (mathmate):

So the old computer can do the job in 19 hours.

OpenStudy (mathmate):

How much of the job does it do in one hour?

OpenStudy (anonymous):

idk @mathmate

OpenStudy (mathmate):

If it takes 19 hours to do the work, in one hour it does 1/19th of the work. Does that make sense?

OpenStudy (anonymous):

yes i get it now @mathmate

OpenStudy (anonymous):

i need help with the last question i sent about the batteries

OpenStudy (mathmate):

Sorry, got to go. Be back afterwards.

OpenStudy (anonymous):

thanks @mathmate

OpenStudy (mathmate):

@brittnicoleee_ Back to the battery problem if you haven't already solved it.

OpenStudy (mathmate):

If there are N batteries in the pack, then the selling price is \[Unit\ price = \frac{47.8}{N}+3.55\] If 12 units recover the cost of the the whole pack, then \[47.8 = 12(\frac{47.8}{N}+3.55)\] Solve for N.

OpenStudy (anonymous):

can you guys help me????????????? Which of the following best describes a chart of the number of tomatoes on a tomato plant each morning? discrete and linear discrete and non-linear continuous and linear continuous and non-linear

OpenStudy (mathmate):

If half or any fraction of a tomato is an acceptable number in the chart, then it is continuous, otherwise it is discrete. You will have to decide whether the number of tomatoes you could get each morning is linear or non-linear.